OpenStudy (anonymous):

Help in Question 2:

OpenStudy (anonymous):

he circuit parameters have the following values: VS=4V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V.

OpenStudy (anonymous):

Given that vIN=0.5V, what is the value of vOUT in Volts? Given that vIN=2V, what is the value of vOUT in Volts? What is the value of vIN in Volts when vOUT drops to VS/2? Now you replace the ExpFET with a MOSFET with parameters K=1mA, Vt = 1v: Again, the circuit parameters are: VS=4V and R=1kΩ, and the ExpFET parameters are K=1mA, VT=1V and VN=1V. Given that VIN=2V, what is the small-signal gain (vout/vin) of this circuit? Given that VIN=2V and assuming that the MOSFET is operating in saturation, what is the small-signal gain (vout/vin) of this circuit?

OpenStudy (anonymous):

p1: Vs p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) put your values in it

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Hey, you, I finished my exam and I got a 60%, but I have a question, wich were the correct answer to problem Q2( b) The circuit parameters have the following values: VS=4V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V. and VIN =2V My answer was 4.39 and was wrong For b) I used v0 = vs-kr(e^((vin-vt)/vn)-1) and I got the wron answer three times, 4.39, The another a,c,d,e were corret using the above formulas except b).